OpenStudy (anonymous):

Express as a single logarithm: 2 log (x-1) + 4 log (x^2+1) - 3 log (x+2)

hartnn (hartnn):

you only need 2 log formulas here : \(\log A+\log B=\log(A+B) \\ \log A^B=B \log A\) try using them....

hartnn (hartnn):

example \(7\log (x+9) =\log (x+9)^7\)

hartnn (hartnn):

so, what about 2 log (x-1) =... ?

OpenStudy (anonymous):

log (x-1)^2

OpenStudy (anonymous):

right?

hartnn (hartnn):

yes, now how about 4 log (x^2+1) =.. ?

hartnn (hartnn):

and 3 log (x+2)= ... ?

hartnn (hartnn):

its the same way you solved 2 log (x-1)

OpenStudy (anonymous):

I'm sorry for being late.

hartnn (hartnn):

no problem,. you can try now :)

OpenStudy (anonymous):

ah! and the answer for my question would be log [(x-1)^2+(x^2+1)^4]/(x+2)^3

hartnn (hartnn):

is there a + sign in between ? there shouldn't be if it looks like : \(\huge \log \frac{(x-1)^2(x^2+1)^4}{(x+2)^3}\) then its correct...

OpenStudy (anonymous):

ah!! thank you! I remembered.. that if there's a +, that means it should be multiplied. right?

hartnn (hartnn):

yes, + sign before log, means multiplication after log.

OpenStudy (anonymous):

oh ok! thanks! I'm really bad at logarithms. Can you help me with another one?

hartnn (hartnn):

sure, but post it as a new question, if i can't help, someone will durely help :)

OpenStudy (anonymous):

wait, it's embarrassing to always post a new question about logarithms. O.o I think you can help me with this.

OpenStudy (anonymous):

Solve for x: \[\log_{2} (x-4) \log_{2} (x-3)=1 \]